Westwind Corporation reports the following results for the current year:

Gross profit on sales                                         $150,000

Long-term capital gain 8000

Long-term capital loss 15000

Western's income tax liability ?

How would your answers to Part a change if Western'sshort-term capital loss is $5,000, instead of $2,000 ?

​short-term capital loss is

$ 5 comma 000$5,000

instead of

$ 2 comma 000$2,000​?

Short-term capital gain 10000

Short-term capital loss 2000

Operating expenses 61000

What are Westwind's taxable income and regular tax liability before credits for the current year?
Answer #1

Part A

Net capital gain = (8000-15000)+(10000-2000) =1000

Gross profit on sales

150000

Plus: Net capital gain

1000

Minus: Operating expenses

(61000)

Taxable income

90000

Tax rate

21%

Tax liability

$18900

Part B

Net capital loss = (8000-15000)+(10000-5000) =$(2000)

Gross profit on sales

150000

Minus: Operating expenses

(61000)

Taxable income

89000

Tax rate

21%

Tax liability

$18690

In any year, if a C corporation's capital losses > capital gains, the excess loss may not be deducted in that year. C corporations are alloed to carryback a net capital for 3 years and carryforward up to a maximum of 5 years.